The Fullname Citation Style

Stuart Shieber

February 2, 1995

The fullname bibliography and citation style is implemented with two files, a BibTEX style file fullname.bst and a L^{A}TEX document-style option file fullname.sty. These are specified in the L^{A}TEX source file as follows:

\documentstyle[fullname]{article}
\bibliographystyle{fullname}

Three citation macros are defined. Normal citations are specified with \cite, which specifies a parenthesized citation giving the authors’ last names (up to three authors) and the year of publication. Here are citations for articles of from one to four authors:

\cite{dewey1} (Dewey, 1910)

\cite{dewey2} (Dewey and Cheatham, 1912)
\cite{dewey3} (Dewey, Cheatham, and Howe, 1914)
\cite{dewey4} (Dewey et al., 1916)

An optional argument allows for specifying further information in the parentheses.

\cite[page 15]{dewey1} (Dewey, 1910, page 15)

\cite[page 15]{dewey2} (Dewey and Cheatham, 1912, page 15)
\cite[page 15]{dewey3} (Dewey, Cheatham, and Howe, 1914, page 15)
\cite[page 15]{dewey4} (Dewey et al., 1916, page 15)

Multiple citations can be placed in a single parenthetical.

\cite{dewey2,cheatham,howe}

(Dewey and Cheatham, 1912; Cheatham, 1918; Howe, 1920)

\cite[page 15]{dewey2,cheatham,howe}

(Dewey and Cheatham, 1912; Cheatham, 1918; Howe, 1920, page 15)

The \namecite macro places the names of the authors outside of and before the parenthesized date. This is useful for citations in the text where the names are part of the sentence, as first described by Dewey and Cheatham (1912).

\namecite{dewey1}  Dewey (1910)

\namecite{dewey2}  Dewey and Cheatham (1912)
\namecite{dewey3}  Dewey, Cheatham, and Howe (1914)

\namecite{dewey4}  Dewey et al. (1916)
\namecite[page 15]{dewey1} Dewey (1910, page 15)

\namecite[page 15]{dewey2} Dewey and Cheatham (1912, page 15)
\namecite[page 15]{dewey3} Dewey, Cheatham, and Howe (1914, page 15)

\namecite[page 15]{dewey4} Dewey et al. (1916, page 15)

The \shortcite macro does not place the authors’ names at all. This is useful if the authors’ names occur in the sentence, but not next to where the citation itself should be. Dewey and Cheatham first used this method (1912). Here are examples of the use of \shortcite in its various forms.

\shortcite{dewey1}  (1910)
\shortcite{dewey2}  (1912)
\shortcite{dewey3}  (1914)
\shortcite{dewey4}  (1916)
\shortcite[page 15]{dewey1} (1910, page 15)
\shortcite[page 15]{dewey2} (1912, page 15)
\shortcite[page 15]{dewey3} (1914, page 15)
\shortcite[page 15]{dewey4} (1916, page 15)

It doesn’t make sense to use \namecite or \shortcite with multiple citations. The output would look like this:

\namecite{dewey2,cheatham,howe}

Dewey and Cheatham (1912; Cheatham (1918; Howe (1920)

\shortcite{dewey2,cheatham,howe}

(1912; 1918; 1920)

The corresponding bibliography style places entries in alphabetical order by first author’s last name, and inverts the first author’s name to emphasize the ordering.

References

Cheatham, Thomas. 1918. The proper style of citation in scholarly papers, and how to spend lots of time getting L^{A}TEX to do it. Life, 12(3):37–38.

Dewey, John. 1910. The proper style of citation in scholarly papers, and how to spend lots of time getting L^{A}TEX to do it. Life, 12(3):37–38.

Dewey, John and Thomas Cheatham. 1912. The proper style of citation in scholarly papers, and how to spend lots of time getting L^{A}TEX to do it. Life, 12(3):37–38.

Dewey, John, Thomas Cheatham, and Elias Howe. 1914. The proper style of citation in scholarly papers, and how to spend lots of time getting L^{A}TEX to do it. Life, 12(3):37–38.

Dewey, John, Thomas Cheatham, Elias Howe, and Stuart Shieber. 1916. The proper style of citation in scholarly papers, and how to spend lots of time getting L^{A}TEX to do it. Life, 12(3):37–38.

Howe, Elias. 1920. The proper style of citation in scholarly papers, and how to spend lots of time getting L^{A}TEX to do it. Life, 12(3):37–38.